Show work please

Complete problem: Required Rate of Return Show your work. Suppose rRF = 6%; rM = 10%; and rA = 14% a. Calculate Stocks A’s beta. b. If Stock A’s beta were 2.0, then what would be A’s new required rate of return?
Answer #1

As per capital asset pricing model of calculating required rate of return of a stock,

Re = Rf + ( Rm – Rf ) x Beta

Where,

Re = Required rate of return = 14%

Rf = Risk free rate of return = 6%

Rm = Market rate of return = 10%

Beta = Beta of the stock

So, putting these values in above equation we get,

14 = 6 + ( 10 – 6 ) x Beta

So, Beta = ( 14 – 6 ) / 4

= 2

So, the beta is 2

If the beta of the stock is 2, the new required rate of return will be

= 6 + ( 10 – 6 ) x 2

= 6 + 8

= 14%